OpenStudy (hocuspucus292333393):

The position of a diving bell is −3/4 mile relative to sea level. The diving bell reached this position after dropping from sea level at a constant rate for  1 1/2 hours. What was the position of the diving bell relative to sea level 1 hour after it began its descent? Enter your answer in the box.

OpenStudy (hocuspucus292333393):

@zoekeya help

OpenStudy (hocuspucus292333393):

@Zeronknight help @OpenStudyRocks5* @Javk @Seratul @Valirium

OpenStudy (hocuspucus292333393):

is it saying it to divide or multiply

OpenStudy (hocuspucus292333393):

):?

OpenStudy (valirium):

@steve816 @ShadowLegendX

OpenStudy (shadowlegendx):

\[-\frac{ 3 }{ 4 }miles = \frac{ 3 }{ 2 }hours\]

OpenStudy (shadowlegendx):

All you have to do is find the factor that converts 3/2hours into 2/2hours, or 1 hour, multiply it by -3/4miles, and you have your answer

OpenStudy (shadowlegendx):

Do so by using the following equation \[\frac{ 3 }{ 2 }x = \frac{ 2 }{ 2 }\] Solve for x Multiply both sides by 2/3 to cancel 3/2 on x (isolating it) \[x = \frac{ 4 }{ 6 } \rightarrow x = \frac{ 2 }{ 3 }\]

OpenStudy (shadowlegendx):

Multiply 3/2hours by 2/3 and you get.... 6/6 hours, or 1 hour. Multiply -3/4 by 2/3 and you get... \[\frac{ -3 }{ 4 } \times \frac{ 2 }{ 3 } = \frac{ -6 }{ 12 } = \frac{ -1 }{ 2 }\]

OpenStudy (shadowlegendx):

For every 1/4 of a mile down, it takes 1/2 of an hour

OpenStudy (hocuspucus292333393):

so the answer is - 1 1/8

OpenStudy (hocuspucus292333393):

um .. hello

OpenStudy (hocuspucus292333393):

\ / - - ___ am I right

OpenStudy (valirium):

Shadowwwww, answer lol

OpenStudy (hocuspucus292333393):

-6/12

OpenStudy (shadowlegendx):

Out of everything I just typed out, all you needed to do was read this "For every 1/4 of a mile down, it takes 1/2 of an hour"

OpenStudy (shadowlegendx):

Though I wished you had been following what I had said, so you could learn ._.

OpenStudy (shadowlegendx):

Look for what the question is asking, and review what I have provided you with.

OpenStudy (hocuspucus292333393):

ok - 6/12 was right?

OpenStudy (danjs):

Maybe you can remember the distance traveled 'd', is equal to the constant rate of traveling 'r' times the time traveled 't' Distance equals rate times time. d = r*t Given the distance -3/4 mi, and the time 3/2 hrs... \[\large \frac{ -3 }{ 4 }=v*\frac{ 3 }{ 2 }\] \[\large v=-\frac{ 1 }{ 2 }\]

OpenStudy (danjs):

The rate the thing is moving is v=-1/2 mi/hr with respect to the sea level. The overall equation is then.. d = (-1/2) * t if they want time t=1 hr, then the distance traveled for that time is. d=-1/2 * 1 = -1/2